OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). Answers y = -4x + 1 y = -1/2 x + 9 y = 4x + 1 y = -2x + 9

OpenStudy (anonymous):

find the slope between the two given points: \(\large m=\frac{y_2-y_1}{x_2-x_1} \) then put it into point-slope form: \(\large y-y_1=m(x-x_1) \) then solve for y to get slope-intercept form....

OpenStudy (anonymous):

try it... what u got for slope? m=

OpenStudy (anonymous):

I don't know I don't understand math at all.

OpenStudy (anonymous):

let's find slope... subtract the y's of your given point... that's what \(\large y_2-y_1 \) means.

OpenStudy (anonymous):

would it be y= -2 +9?

OpenStudy (anonymous):

is that a guess or did you actually do what i said?

OpenStudy (anonymous):

I actually did what you said.